Using the Intermediate Value Theorem and a calculator, find an interval of length 0.01 that contains a root of e^x = 2 - x, rounding interval endpoints off to the nearest hundredth.


1
Expert's answer
2021-10-25T17:18:46-0400

The function "g(x)=e^x" is strictly increasing. The function "h(x)=2-x" is strictlydecreasing.

Then if the root of "e^x=2-x" exists, it will be the only root.

The function"f(x)=e^x-(2-x)" is continuous on "(-\infin, \infin)."


"f(0)=e^{0}-(2-0)=-1<0"

"f(0.5)=e^{0.5}-(2-0.5)=e^{0.5}-1.5=\sqrt{e}-1.5>0"

"f(0.2)=e^{0.2}-(2-0.2)=e^{0.2}-1.8\approx-0.5789<0"

"f(0.4)=e^{0.4}-(2-0.4)=e^{0.4}-1.6\approx-0.108<0"

By the Intermediate Value Theorem exists a number "c\in (0.4,0.5)" such that "f(c)=0"


"f(0.45)=e^{0.45}-(2-0.45)=e^{0.45}-1.55\approx0.018>0"

"f(0.42)=e^{0.42}-(2-0.42)=e^{0.42}-1.58\approx-0.058<0"

"f(0.44)=e^{0.44}-(2-0.44)=e^{0.44}-1.56\approx-0.007<0"

By the Intermediate Value Theorem exists a number "c\in (0.44,0.45)," such that "f(c)=0."


"(0.44,0.45)"